# The Optimal Strategy for Playing Lucky 13

### v1

@@ Abstract
The game show Lucky 13 differs from other television game shows in that contestants are required to place a bet on their own knowledge of trivia by selecting a range that contains the number of questions that they answered correctly. We present a model for this game show using binomial random variables and generate tables outlining the optimal range the player should select based on maximization of two different utility functions. After analyzing the decisions made by some actual contestants on this show, we present a numerical simulation for how many questions an average player is expected to answer correctly based on question categories observed for two sample contestants.
@@ 1 Introduction
Strategic decision making in games provides a unique and entertaining application of probability theory. There are strategies available for most common casino games. For example, in [1], the optimal strategy for blackjack is discussed depending on whether or not the player wants to count cards. Many television game shows also provide scenarios in which strategic play must be applied. Perhaps the most famous example is the Monty Hall Problem commonly seen in almost every introductory probability book. The original version of the game and some of its variations are discussed in [2]. In television game shows where contestants are asked to answer trivia questions, one can usually win more money by answering more questions correctly. For example, in the popular game show “Who wants to be a millionaire” contestants were asked a series of 14 or 15 questions and were allowed to use a set of lifelines on questions or walk away from a question if they were not confident. Using stochastic control techniques, Dalang and Bernyk [3] have created tables to show the optimal lifeline strategy, upper bounds on the expected payoff when the optimal strategy is followed, and the upper bounds on the probability of winning the million dollars given a certain number of questions have been correctly answered. However, in the 2024 ABC television game show Lucky 13, contestants not only had to answer the questions correctly, but also correctly predict how many correct answers they gave. In this scenario, your ability to evaluate your own knowledge is just as important as providing the maximum number of correct responses.
@@ 2 Game Description
Before playing this game, a contestant probably has already decided how much value each cash prize has. For example, a wealthy businessman who regularly plays trivia games would probably place a higher personal value on winning \$1,000,000 than on winning any lower prize. However, other contestants may simply be interested in winning some amount of money and would be less willing to engage in risky play. Mathematically, this personal value of the cash prizes can be quantified by a utility function that maps each prize to a real number where higher values correspond to a more desirable personal value. Appropriate mathematical axioms needed to define a utility function on a discrete set are given in [4]. In this paper, we apply our models to two different utility functions. Our models can be applied to other utility functions should a player choose a different one than the ones discussed here.
@@ 3 Model with Two Question Categories
In order to determine a strategy for correctly selecting the Lucky Range and Lucky Number, we consider a model where each of the 13 questions can be categorized into one of two different types. A questions is “Sure” (S) if the contestant gives the correct answer with probability 1. A question is “Guess” (G) if the contestant will give the correct answer with probability 0.5. Let $N_{S}$ be the number of Sure questions (out of 13) a contestant has. Let $N_{G}$ be the number of Guess questions that the contestant answers correctly. These are the extreme cases since we assume either a contestant knows the answer for sure or randomly guesses. We also assume that the contestant is able to decide which category is appropriate for each question. In the next section we introduce a model that takes partial knowledge of a question into account. Clearly, $N_{S}$ is not random but $N_{G}$ is a random variable that follows a binomial distribution. More specifically, $N_{G}\sim\text{Binom}(13-N_{S},0.5)$. Additionally, define the random variable $N$ to be the total number of correct answers out of 13, so that $N=N_{S}+N_{G}$.
@@
Table 3: Table of Lucky Ranges and Expected Winnings when $N_{1}=3$
@@
In this case, Contestant $Y$ will need to apply some personal judgment in order to select his Lucky Range. If he wants to maximize the probability he wins some money, he should choose Lucky Range 7-9 because this one has the greatest probability. However, we see from Table 3 that the mathematical expectation for his winnings is higher if he picks Lucky Range 10-12. Although it is not possible for Contestant $Y$ to win exactly \$17,090 (the mathematical expected winning), selecting Lucky Range 10-12 would be the better strategy in the long term if Contestant $Y$ were playing this game multiple times and always was sure about 3 answers. However, for the single time he must make his decision, he needs to evaluate how much the cash prizes are worth to him. If Contestant $Y$ only wants to win a large sum of money and likes to take risks, he should select 10-12. On the other hand, if he is a more risk-adverse player, he should select 7-9. This example shows that for this game show, maximizing the probability of winning does not necessarily lead to selecting the same Lucky Range as a player who chooses to maximize his potential cash winnings.
@@ 4 Model with Three Question Categories
${2\choose 0}\left(\frac{1}{4}\right)^{2}\cdot{1\choose 1}\left(\frac{1}{2}\right)^{1}+{2\choose 2}\left(\frac{3}{4}\right)^{2}\cdot{1\choose 0}\left(\frac{1}{2}\right)^{1}=\frac{7}{32}$
@@
There are 105 different combinations of Sure, Unsure, and Guess questions. We present the optimal decision based on the two different utility functions for choosing a Lucky Number and Lucky Range for selected combinations in Table 6. The strategy for all combinations of $S,U,$ and $G$ can be found in Appendix A. We see that it is never optimal to select Lucky Range 1-3 and risk-adverse players should only select Lucky Range 4-6 if they guess on all 13 questions. Maximizing expected winnings corresponds to more risky play when a contestant’s expected number of correct answers is between one of the Lucky Ranges in which case they choose the range with higher cash prize (even if there is a lower probability of being in that range).
@@ 8 Simulation of Number of Correct Responses
We assume that a player that is randomly selected has an expertise in a random number of these categories. If the player has an expertise in the question category, we assume that question will be $U$ with probability 0.6 and $S$ with probability 0.4 which would make the total probability the player gives a correct answer 0.85. Otherwise, we assume the player will supply the correct answer with probability 0.5. The number of categories a randomly selected player has an expertise on is modeled by $Q:=R+1$ where $R\sim\text{Binom}(20,1.5/20)$. The parameter of 20 was selected since there were 20 categories of questions in Table 7 and the success probability was chosen so that on average a player has expertise in 2.5 categories. We sample without replacement from the set of question categories given in Table 7 in order to select which categories a player will have an expertise in. For each of the 10,000 players in our Monte Carlo simulation we compute $Q$, take a sample of size $Q$ from the set of all categories, and have the player answer 13 randomly selected questions with the modified probabilities discussed above if a player has an expertise in a category. The distribution of correct responses under this model is given in Figure 6. We observe that the most frequently occurring numbers of correct responses are 6,7, or 8. Thus, it would be wise for Lucky 13 contestants to memorize the correct Lucky Number based on their utility function for values $S,U,$ and $G$ such that $6\leq\text{E}[N]\leq 8$. We also note that it is possible to determine the exact distribution of the number of correct answers but we chose to do a Monte Carlo simulation instead.
@@ 9 Model with Zero Question Categories
The most general model for this game is one where the contestant may specify $\vec{p}=(p_{1},p_{2},\cdots,p_{13})$ where $p_{i}$ is the probability you get the ith question correct. If we again define $N$ as the total number of correct responses a contestant gives, $N$ has a Poisson binomial distribution with expected value $\mu=\sum_{i=1}^{13}p_{i}$. See [5] for more details on the Poisson binomial distribution. Note that instead of categories of questions, this model assigns any probability $0.5\leq p_{i}\leq 1$ to each question. Tang and Tang (see [5]) give Darroch’s rule for the mode ($m$) of a Poisson binomial random variable based on its mean $\mu$:

### v2

@@ Abstract
The game show Lucky 13 differs from other television game shows in that contestants are required to place a bet on their own knowledge of trivia by selecting a range that contains the number of questions that they answered correctly. We present several models for this game show using binomial random variables and generate tables outlining the optimal range the player should select based on maximization of two different utility functions. After analyzing the decisions made by some actual contestants on this show, we present a numerical simulation for how many questions an average player is expected to answer correctly based on question categories observed for twelve sample contestants.
@@ 1 Introduction
Strategic decision making in games provides a unique and entertaining application of probability theory. There are strategies available for most common casino games. For example, in [Bib3], the optimal strategy for blackjack is discussed depending on whether or not the player wants to count cards. Many television game shows also provide scenarios in which strategic play must be applied. Perhaps the most famous example is the Monty Hall Problem commonly seen in almost every introductory probability book. The original version of the game and some of its variations are discussed in [Bib4]. In television game shows where contestants are asked to answer trivia questions, one can usually win more money by answering more questions correctly. For example, in the popular game show “Who wants to be a millionaire” contestants were asked a series of 14 or 15 questions and were allowed to use a set of lifelines on questions or walk away from a question if they were not confident. Using stochastic control techniques, Dalang and Bernyk [Bib1] have created tables to show the optimal lifeline strategy, upper bounds on the expected payoff when the optimal strategy is followed, and the upper bounds on the probability of winning the million dollars given a certain number of questions have been correctly answered. However, in the 2024 ABC television game show Lucky 13, contestants not only had to answer the questions correctly, but also correctly predict how many correct answers they gave. In this scenario, your ability to evaluate your own knowledge is just as important as providing the maximum number of correct responses.
@@ 2 Game Description
Before playing this game, a contestant probably has already decided how much value each cash prize has. For example, a wealthy businessman who regularly plays trivia games would probably place a higher personal value on winning \$1,000,000 than on winning any lower prize. However, other contestants may simply be interested in winning some amount of money and would be less willing to engage in risky play. Mathematically, this personal value of the cash prizes can be quantified by a utility function that maps each prize to a real number where higher values correspond to a more desirable personal value. Appropriate mathematical axioms needed to define a utility function on a discrete set are given in [Bib2]. In this paper, we apply our models to two different utility functions. Our models can be applied to other utility functions should a player choose a different one than the ones discussed here.
@@ 3 Model with Two Question Categories
In order to determine a strategy for correctly selecting the Lucky Range and Lucky Number, we consider a model where each of the 13 questions can be categorized into one of two different types. A question is “Sure” (S) if the contestant gives the correct answer with probability 1. A question is “Guess” (G) if the contestant will give the correct answer with probability 0.5. Let $N_{S}$ be the number of Sure questions (out of 13) a contestant has. Let $N_{G}$ be the number of Guess questions that the contestant answers correctly. These are the extreme cases since we assume either a contestant knows the answer for sure or randomly guesses. We also assume that the contestant is able to decide which category is appropriate for each question. In the next section we introduce a model that takes partial knowledge of a question into account. Clearly, $N_{S}$ is not random but $N_{G}$ is a random variable that follows a binomial distribution. More specifically, $N_{G}\sim\text{Binom}(13-N_{S},0.5)$. Additionally, define the random variable $N$ to be the total number of correct answers out of 13, so that $N=N_{S}+N_{G}$.
@@
Table 3: Table of Lucky Ranges and Expected Winnings when $N_{S}=3$
@@
In this case, Contestant $Y$ will need to apply some personal judgment in order to select his Lucky Range. If he wants to maximize the probability he wins some money, he should choose Lucky Range 7-9 because this one has the greatest probability. However, we see from Table 3 that the mathematical expectation for his winnings is higher if he picks Lucky Range 10-12. Although it is not possible for Contestant $Y$ to win exactly \$17,090 (the mathematical expected winning), selecting Lucky Range 10-12 would be the better strategy in the long term if Contestant $Y$ were playing this game multiple times and always was sure about 3 answers. However, for the single time he must make his decision, he needs to evaluate how much the cash prizes are worth to him. If Contestant $Y$ only wants to win a large sum of money and likes to take risks, he should select 10-12. On the other hand, if he is a more risk-averse player, he should select 7-9. This example shows that for this game show, maximizing the probability of winning does not necessarily lead to selecting the same Lucky Range as a player who chooses to maximize his potential cash winnings.
@@ 4 Model with Three Question Categories
${2\choose 0}\left(\frac{1}{4}\right)^{2}\cdot{1\choose 1}\left(\frac{1}{2}\right)^{1}+{2\choose 1}\left(\frac{3}{4}\right)^{1}\left(\frac{1}{4}\right)^{1}\cdot{1\choose 0}\left(\frac{1}{2}\right)^{1}=\frac{7}{32}$
@@
There are 105 different combinations of Sure, Unsure, and Guess questions. We present the optimal decision based on the two different utility functions for choosing a Lucky Number and Lucky Range for selected combinations in Table 6. The strategy for all combinations of $S,U,$ and $G$ can be found in Appendix A. We see that it is never optimal to select Lucky Range 1-3 and risk-averse players should only select Lucky Range 4-6 if they guess on all 13 questions. Maximizing expected winnings corresponds to more risky play when a contestant’s expected number of correct answers is between one of the Lucky Ranges in which case they choose the range with higher cash prize (even if there is a lower probability of being in that range).
@@ 8 Simulation of Number of Correct Responses
We assume that a player that is randomly selected has an expertise in a random number of these categories. If the player has an expertise in the question category, we assume that question will be $U$ with probability 0.6 and $S$ with probability 0.4 which would make the total probability the player gives a correct answer 0.85. Otherwise, we assume the player will supply the correct answer with probability 0.5. The number of categories a randomly selected player has an expertise on is modeled by $Q:=R+1$ where $R\sim\text{Binom}(20,1.5/20)$. The parameter of 20 was selected since there were 20 categories (excluding the “Other” category) of questions in Table 7 and the success probability was chosen so that on average a player has expertise in 2.5 categories. We sample without replacement from the set of question categories given in Table 7 in order to select which categories a player will have an expertise in. For each of the 10,000 players in our Monte Carlo simulation we compute $Q$, take a sample of size $Q$ from the set of all categories, and have the player answer 13 randomly selected questions with the modified probabilities discussed above if a player has an expertise in a category. The distribution of correct responses under this model is given in Figure 6. We observe that the most frequently occurring numbers of correct responses are 6,7, or 8. Thus, it would be wise for Lucky 13 contestants to memorize the correct Lucky Number based on their utility function for values $S,U,$ and $G$ such that $6\leq\text{E}[N]\leq 8$. We also note that it is possible to determine the exact distribution of the number of correct answers but we chose to do a Monte Carlo simulation instead.
@@
9 Extended Three Category Model
The major flaw with the previously discussed three-category model is that contestants may have significant variability in the estimation of the probability that they answer an Unsure question correctly. In order to study this scenario, we will leave the Sure and Guess categories unchanged and modify the model for the Unsure category. Now, for all questions that are not Sure or Guess, we will assign each of them probability $\pi$ of being a Sure question and probability $1-\pi$ of being a Guess question. We can then impose a prior probability distribution on $\pi$ which leads us to the following theorem.
Theorem 1.
If $X_{i}\sim\text{Bern}(\pi_{i})$ and $\pi_{i}\stackrel{{\scriptstyle iid}}{{\sim}}F$ with CDF $F(\pi)$ being any probability distribution on $[0,1]$ with PDF $f(\pi)$, then the marginal distribution of $X_{i}$ is $\text{Bern}(\text{E}[\pi_{i}])$.
Proof.
Let $N_{U}$ be the number of questions answered correctly out of $U$ Unsure questions. Clearly, $N_{U}=\sum_{i=1}^{U}X_{i}$ where $X_{i}$ is the outcome of the $i$-th question. Let $\pi_{i}$ be an independent sample from $f(\pi)$. Thus,
$\text{P}(X_{i}=1)=\int_{0}^{1}\text{P}(X_{i}=1|\pi_{i})f(\pi_{i})d\pi_{i}$
$\int_{0}^{1}\pi_{i}f(\pi_{i})d\pi_{i}$
$\text{E}[\pi].$
Since $\pi_{i}$ are independent, $X_{i}$ are marginally independent Bernoulli trials with success probability $\text{E}[\pi]$. Thus, $N_{U}\sim\text{Binom}(U,\text{E}[\pi])$. ∎
Corollary 2.
Any two prior distributions with the same mean will yield the same strategy.
Proof.
Since the values of $N_{S}$ and $N_{G}$ did not depend on $\pi_{i}$, any two prior distributions with the same mean will yield the same values for $\text{P}(N=k|\vec{\pi})$ where $k\in[0,13]$ and $\vec{\pi}=(\pi_{1},\cdots,\pi_{U})$. ∎
The uniform distribution is a natural choice for the prior distribution of $\pi_{i}$. For this choice of prior, we can easily calculate the expected value and variance of
$N=S+\text{Binom}\left(G,\frac{1}{2}\right)+\sum_{i=1}^{U}\text{Bern}\left(\pi_{i}\cdot 1+(1-\pi_{i})\cdot\frac{1}{2}\right)$
$S+\text{Binom}\left(G,\frac{1}{2}\right)+\sum_{i=1}^{U}\text{Bern}\left(\frac{\pi_{i}+1}{2}\right).$
Assuming $\pi_{i}\sim\text{Unif}[a_{i},b_{i}]$ for real numbers $a_{i},b_{i}$ such that $0\leq a_{i}\leq b_{i}\leq 1$, we find that
$\text{E}[N]=\text{E}[\text{E}[N|\vec{\pi}]]$
$\text{E}\left[S+\frac{G}{2}+\sum_{i=1}^{U}\text{E}\left[\text{Bern}\left(\frac{\pi_{i}+1}{2}\right)\right]\right]$
$S+\frac{G}{2}+\sum_{i=1}^{U}\text{E}\left[\frac{\pi_{i}+1}{2}\right]$
$S+\frac{G}{2}+\sum_{i=1}^{U}\left(\frac{1}{2}\cdot\frac{a_{i}+b_{i}}{2}+\frac{1}{2}\right)$
$S+\frac{G}{2}+\frac{U}{2}+\frac{1}{4}\sum_{i=1}^{U}(a_{i}+b_{i})$
and
$\text{Var}[N]=\text{E}[\text{Var}[N|\vec{\pi}]]+\text{Var}[\text{E}[N|\vec{\pi}]]$
$\text{E}\left[0+\frac{G}{4}+\sum_{i=1}^{U}\text{Var}\left[\text{Bern}\left(\frac{\pi_{i}+1}{2}\right)\right]\right]+\text{Var}\left[S+\frac{G}{2}+\sum_{i=1}^{U}\left(\frac{\pi_{i}+1}{2}\right)\right]$
$\frac{G}{4}+\sum_{i=1}^{U}\text{E}\left[\left(\frac{\pi_{i}}{2}+\frac{1}{2}\right)\left(\frac{1}{2}-\frac{\pi_{i}}{2}\right)\right]+\frac{1}{4}\sum_{i=1}^{U}\text{Var}[\pi_{i}]$
$\frac{G}{4}+\frac{1}{4}\sum_{i=1}^{U}\text{E}[1-\pi_{i}^{2}]+\frac{1}{48}\sum_{i=1}^{U}(b_{i}-a_{i})^{2}$
$\frac{G}{4}+\frac{U}{4}-\frac{1}{4}\sum_{i=1}^{U}\left(\frac{(b_{i}-a_{i})^{2}}{12}+\frac{(a_{i}+b_{i})^{2}}{4}\right)+\frac{1}{48}\sum_{i=1}^{U}(b_{i}-a_{i})^{2}$
$\frac{G}{4}+\frac{U}{4}-\frac{1}{16}\sum_{i=1}^{U}(a_{i}+b_{i})^{2}.$
We now demonstrate the impact of $\vec{\pi}$ on a Lucky Number decision. Assume a player has $S=9,U=2,\text{ and }G=2$. Assuming we are also provided with the vector $(\pi_{1},\pi_{2})$, we can compute the following probabilities:
$\text{P}(N=13|\vec{\pi})=\text{P}(N_{U}=2|\vec{\pi})\text{P}(N_{G}=2)$
$\left(\frac{\pi_{1}+1}{2}\right)\left(\frac{\pi_{2}+1}{2}\right){2\choose 0}\left(\frac{1}{2}\right)^{2}$
$\frac{1}{16}[\pi_{1}\pi_{2}+\pi_{1}+\pi_{2}+1]$
$\text{P}(N=12|\vec{\pi})=\text{P}(N_{U}=2|\vec{\pi})\text{P}(N_{G}=1)+\text{P}(N_{U}=1|\vec{\pi})\text{P}(N_{G}=2)$
$\left(\frac{\pi_{1}+1}{2}\right)\left(\frac{\pi_{2}+1}{2}\right){2\choose 1}\left(\frac{1}{2}\right)^{2}+\left(\frac{\pi_{1}+1}{2}\right)\left(\frac{1-\pi_{2}}{2}\right){2\choose 2}\left(\frac{1}{2}\right)^{2}+\left(\frac{1-\pi_{1}}{2}\right)\left(\frac{\pi_{2}+1}{2}\right)\left(\frac{1}{2}\right)^{2}$
$\frac{1}{16}[2\pi_{1}+2\pi_{2}+4]$
$\text{P}(N=11|\vec{\pi})=\text{P}(N_{U}=2|\vec{\pi})\text{P}(N_{G}=0)+\text{P}(N_{U}=0|\vec{\pi})\text{P}(N_{G}=2)+\text{P}(N_{U}=1|\vec{\pi})\text{P}(N_{G}=1)$
$\frac{1}{16}[6-2\pi_{1}\pi_{2}]$
$\text{P}(N=10|\vec{\pi})=\text{P}(N_{U}=1|\vec{\pi})\text{P}(N_{G}=0)+\text{P}(N_{U}=0|\vec{\pi})\text{P}(N_{G}=1)$
$\frac{1}{16}[4-2\pi_{2}-2\pi_{1}]$
$\text{P}(N=9|\vec{\pi})=\text{P}(N_{U}=0|\vec{\pi})\text{P}(N_{G}=0)$
$\frac{1}{16}[1-\pi_{1}-\pi_{2}+\pi_{1}\pi_{2}].$
Now, we can investigate which Lucky Range and Lucky Number this player should choose in order to maximize the probability of winning. In order for 13 to have maximum probability over Lucky Range 10-12, we would need to have:
$\text{P}(N=13|\vec{\pi})>\sum_{i=10}^{12}\text{P}(N=i|\vec{\pi})$
which is equivalent to
$\pi_{1}+\pi_{2}+\pi_{1}\pi_{2}+1>2\pi_{1}+2\pi_{2}+4+6-2\pi_{1}\pi_{2}+4-2\pi_{2}-2\pi_{1}$
or
$\pi_{1}+\pi_{2}+3\pi_{1}\pi_{2}>13$
which clearly cannot happen since $\pi_{1},\pi_{2}\in[0,1]$. Thus, claiming all questions were answered correctly is not a better strategy than selecting Lucky Range 10-12. Now we compare Lucky Ranges 10-12 and 7-9. In order for Lucky Range 10-12 to have the maximum probability, we would need
$\sum_{i=10}^{12}\text{P}(N=i|\vec{\pi})>\text{P}(N=9|\vec{\pi})$
which is equivalent to
$2\pi_{1}+2\pi_{2}+4+6-2\pi_{1}\pi_{2}+4-2\pi_{2}-2\pi_{1}>1-\pi_{2}-\pi_{1}+\pi_{1}\pi_{2}$
or
$13>3\pi_{1}\pi_{2}-\pi_{2}-\pi_{1}.$
This is always true. In other words, to maximize the probability of winning, a contestant with $S=9,U=2,\text{ and }G=2$ should always select Lucky Range 10-12. Now we compare the probabilities of each Lucky Number in the range 10-12. Note that $P_{10}:=\text{P}(N=10|\vec{\pi})-4=-2\pi_{2}-2\pi_{1}$ and $P_{11}:=\text{P}(N=11|\vec{\pi})-4=2(1-\pi_{1}\pi_{2})$ and $P_{12}:=\text{P}(N=12|\vec{\pi})-4=2(\pi_{1}+\pi_{2})$. It is clear that $P_{10}<\min\{P_{11},P_{12}\}$ so we only need to compare $P_{11}$ and $P_{12}$. If $1-\pi_{1}\pi_{2}>\pi_{1}+\pi_{2}$ we will select Lucky Number 11. However, this condition is equivalent to $(1-\pi_{1})-\pi_{2}(1+\pi_{1})>0$ which implies $\pi_{2}<\frac{1-\pi_{1}}{1+\pi_{1}}$. Figure 7 shows the regions where we pick Lucky Number 11 versus Lucky Number 12.
Figure 7: Impact of $\vec{\pi}$ when $S=9,U=2,\text{ and }G=2$
The area of the region shaded in Figure 7 is given by
$\int_{0}^{1}\frac{1-\pi_{1}}{1+\pi_{1}}d\pi_{1}=2\ln(2)-1\approx 0.38.$
This probability represents the proportion of times that a player would select 11 assuming that the prior distribution of $\pi_{1}$ and $\pi_{2}$ are Unif[0,1]. Additionally, Table 9 shows the corresponding values of $S$, $U$, and $G$, for various points in Figure 7.
Table 8: Table of Coordinate Point Equivalencies for $S=9,U=2,\text{ and }G=2$
Point $S/U/G$
$(0,0)$ 9/0/4
$(1,0)$ 10/0/3
$(0,1)$ 10/0/3
$(1,1)$ 11/0/2
[…]
@@ 10 Model with Zero Question Categories
The most general model for this game is one where the contestant may specify $\vec{p}=(p_{1},p_{2},\cdots,p_{13})$ where $p_{i}$ is the probability you get the ith question correct. If we again define $N$ as the total number of correct responses a contestant gives, $N$ has a Poisson binomial distribution with expected value $\mu=\sum_{i=1}^{13}p_{i}$. See [Bib5] for more details on the Poisson binomial distribution. Note that instead of categories of questions, this model assigns any probability $0.5\leq p_{i}\leq 1$ to each question. Tang and Tang (see [Bib5]) give Darroch’s rule for the mode ($m$) of a Poisson binomial random variable based on its mean $\mu$: